# MaskINT: Video Editing via Interpolative Non-autoregressive Masked Transformers

### v1

@@ 4.3 MaskINT
Previous works [16, 51] demonstrate that non-autoregressive masked generative transformers can effectively perform frame prediction and interpolation tasks. However, these architectures lack explicit control of structure, making it difficult to follow the motion of original videos. In our work, we explicitly introduce structural condition of each frame into the generation process. Specifically, we tokenize both RGB frames $\{\mathbf{I}^{i}\}_{i=1}^{N}$ and structure maps $\{\mathbf{S}^{i}\}_{i=1}^{N}$ with an off-the-shelf 2D VQ tokenizer [12]. We utilize 2D VQ rather than 3D VQ [51] to accommodate varying numbers of frames and frame rate without constraints. We denote the tokens from RGB frames as $\mathbf{{}^{c}Z}=\{^{c}z_{i}\}_{i=1}^{N\times h\times w}$ (color token) and tokens from edge maps as $\mathbf{{}^{s}Z}=\{^{s}z_{i}\}_{i=1}^{N\times h\times w}$ (structure token), where ${}^{s}z_{i}$, ${}^{c}z_{i}\in\{1,2,...,M\}$, where $M$ is the codebook size. Subsequently, two distinct embedding layers $e^{C}(\cdot)$ and $e^{S}(\cdot)$ are learned to map token indices $\mathbf{{}^{c}Z}$ and $\mathbf{{}^{s}Z}$ into their respective embedding spaces. Learnable 2D spatial positional encoding $\mathbf{P^{S}}$ and temporal positional encoding $\mathbf{P^{T}}$ are also added [2]. Thus, the input to the following transformer layers can be formulated by $\mathbf{X}=e^{c}(\mathbf{{}^{c}Z})+e^{s}(\mathbf{{}^{s}Z})+\mathbf{P^{S}}+\mathbf{P^{T}}\in\mathbb{R}^{N\times h\times w\times c}$.
@@
Previous masked generative transformers [7, 51] employ a pure transformer with global attention [44]. However, Given that there is no substantial motion between consecutive frames, we adopt self-attention within a restricted window to further mitigate computational overhead, following MaskViT [16]. In detail, our approach involves two distinct stages of attention. Initially, we employ spatial window attention, confining attention to tokens within a frame of dimensions $1\times h\times w$. Subsequently, we extend this to spatial-temporal window attention, which confines attention to tokens within a tube of dimensions $N\times h_{w}\times w_{w}$, where $h_{w}$ and $w_{w}$ is the window size, where $h_{w}<h$ and $w_{w}<w$, which greatly reduce the complexity. Besides, to further reduce computation of transformer, we also add a shallow convolution layers to downsample $\mathbf{X}$ before the transformer encoder layers and an upsample layer at the end.
@@
By fully disentangling keyframes editing and frame interpolation into distinct stages, our model no longer necessitates paired videos for training. Consequently, we can train MaskINT using video only datasets. Denote the color token of $i^{th}$ RGB frame as $\mathbf{{}^{c}Z}^{i}=\{^{c}z^{i}\}_{i=1}^{h\times w}$. During the training time, we keep color tokens of the initial frame $\mathbf{{}^{c}Z}^{1}$ and last frame $\mathbf{{}^{c}Z}^{N}$, and randomly replace $[\gamma(r)\cdot(N-2)\cdot N]$ color tokens of intermediate frames with the [MASK] tokens. We denote this corrupted video color tokens as $\mathbf{\bar{{}^{c}Z}}=\{\mathbf{{}^{c}Z}^{1},\mathbf{\bar{{}^{c}Z}}^{2},...,\mathbf{\bar{{}^{c}Z}}^{N-1},\mathbf{{}^{c}Z}^{N}\}$. The structure-aware window-restricted transformer with parameters $\Theta$ is trained by
$\mathcal{L}_{MTM}=\mathop{\mathbb{E}}_{\mathbf{{}^{c}Z},\mathbf{{}^{s}Z}\in\mathbb{D}}[\sum_{\bar{{}^{c}z_{i}}=\texttt{[MASK]}}-\log p_{\Theta}(^{c}z_{i}|\mathbf{\bar{{}^{c}Z}},\mathbf{{}^{s}Z})]$ (4)
@@
During the inference time, our MaskINT can seamlessly generalize to perform frame interpolation between the jointly edited frames, although it is only trained with regular videos. Specifically, we tokenize the the initial and last edited frames $\mathbf{{}^{*}I}^{1}$ and $\mathbf{{}^{*}I}^{N}$ from Stage 1 into color tokens $\mathbf{{}^{c}_{*}Z}^{1}$ and $\mathbf{{}^{c}_{*}Z}^{N}$, and initialize color tokens of all intermediate frames $\{\mathbf{\bar{{}^{c}Z}}^{2},...,\mathbf{\bar{{}^{c}Z}}^{N-1}\}$ with [MASK] tokens. We follow the iterative decoding in MaskGiT [7] with a total number of $K$ steps. At step $k$, we predict all color tokens in parallel and keep tokens with the highest confidence score.
@@ 5.1 Settings
We train our model with $100k$ videos from ShutterStock website [^1] https://www.shutterstock.com/video. During training time, we random select a $T=16$ video clip with frame interval $1,2,4$ from each video and resize it to $384\times 672$. We utilize Retina-VQ [11] with $8$ downsample ratio, i.e., each frame has $48\times 84$ tokens. We employ Transformer-Base as our MaskINT and optimized it from scratch with the AdamW optimizer [27] for a duration of $100$ epochs. The initial learning rate is set to $41e-$ 4 and decayed with cosine schedule. During the inference time, we set the number of decoding step $K$ to $32$ and the temperature $t$ to $4.5$.
@@
Following [47], we use the selected $40$ object-centric videos of the DAVIS dataset [33], covering humans, animals, vehicles, etc. Besides, we also select $30$ unseen videos from the ShutterStock dataset. For each video, we manually design $5$ edited prompts, including object editing, background changes and style transfers. Following previous works [53, 34], we assess the quality of the generated videos using CLIP [35]. In detail, we evaluate 1) Temporal consistency, which calculates the average cosine similarity of all pairs of consecutive frames. 2) Prompt consistency, which calculates the average cosine similarity between given text prompt and all video frames. To evaluate the efficiency, we report the duration required for generating a 16-frame video clip on a single NVIDIA A6000 GPU.
@@ 5.2 Results
We select methods that built upon text-to-image diffusion models for comparison, including TokenFlow [15], Text-to-video zero [25], and ControlVideo [53]. We also consider apply ControlNet to each frame individually with the same initial noise as baseline. Besides, we also compare with frame interpolation with FILM [36] with the same edited keyframes.
@@
T.C. P.C. T.C. P.C
ControlNet per frame [52] 0.9137 0.3136 0.9419 0.3040 50s
Text2Video-zero [25] 0.9642 0.3124 0.9811 0.3036 60s
TokenFlow [15] 0.9774 0.3169 0.9869 0.3133 150s
ControlVideo-edge [53] 0.9746 0.3143 0.9864 0.3032 120s
MaskINT (ours) 0.9519 0.3112 0.9714 0.3038 22s
Table 1: Quantitative comparisons on the quality of the generated videos. “T.C.” stands for “temporal consistency”, and “P.C.” stands for “prompt consistency”.
@@
Table 1 summarize the performance of these methods on both DAVIS and ShutterStock datasets. Notably, our method achieves comparable performance with diffusion methods, in terms of both temporal consistency and prompt consistency, while brings a significant acceleration in processing speed. In detail, MaskINT is almost 5.5 times faster than ControlVideo [53], whose fully cross-frame attention is computationally extensive. Moreover, MaskINT is nearly 7 times faster than TokenFlow [15], whose DDIM inversion is time-consuming. On the contrary, our acceleration is derived from a combination of a lightweight network design and a reduced number of decoding steps in masked generative transformers.
@@
Fig. 1 and Fig. 3 show several samples of edited videos with MaskINT. Our method is capable of generating temporally consistent videos that adhere to the provided text prompts. This extends to a wide range of applications with text prompts, encompassing tasks such as stylization, background editing, foreground editing, and more. It also works well on challenging videos with substantial motion, such as jumping and running. Moreover, Fig. 4 provides a qualitative comparison of MaskINT to other baselines. Remarkably, diffusion methods [15, 25, 53] can ensure the consistency of overall appearance, but sometimes cannot maintain the consistency of detailed regions. For example, both TokenFlow [15] and Text2Video-Zero[25] exhibit noticeable artifacts in the leg region of the human subjects. ControlVideo [53] produces inconsistent hats. The potential explanation lies in the fact that these methods offer control over temporal consistency implicitly. Furthermore, FILM [36] produces videos that deviate from the original motions with the same edited keyframes. Our MaskINT consistently interpolates the intermediate frames based on the structure condition and even maintain better consistency in local regions.
@@
Since the non-autoregressive pipeline generates all video frames simultaneously, it’s challenging for it to edit an entire long video due to GPU memory limitation. Nevertheless, our framework can still be extended to generate long videos by dividing the long video into short clips and progressively performing frame interpolation within each clip. For instance, given a video with $60$ frames, we select the $1^{\text{st}}$, $16^{\text{th}}$, $31^{\text{st}}$, $46^{\text{th}}$, and $60^{\text{th}}$ frames as keyframes. We jointly edit these selected $5$ frames together and then perform structure-aware frame interpolation within each pair of consecutive keyframes. As shown in Fig. 5, with this design, our method can still generate consistent long videos. Besides, in this proposed extension, the generation of later frames is decoupled from the generated early frames, which differs from the autoregressive long-video generation pipeline in Phenaki [45]. Consequently, even if some early frames encounter difficulties, the generation of later frames can still proceed successfully.
@@ 5.3 Ablation Studies
Video Frame Interpolations
We conduct a quantitative evaluation on the performance of our structure-aware interpolation module. In this ablation study, we perform frame interpolation using the original keyframes to reconstruct the original intermediate frames, and compare the interpolated frames with the original video frames at the pixel level. We use same testing samples from DAVIS and Shutterstock datasets, employing peak signal-to-noise ratio (PSNR), learned perceptual image patch similarity (LPIPS), and structured similarity (SSIM) as the evaluation metrics. We benchmark our method against two state-of-the-art VFI methods, namely FILM [36] and RIFE [22]. We also showcase the performance of applying VQ-GAN [12] to all video frames, serving as an upper bound for our method.
As in Table 2, our method significantly outperforms VFI methods on all evaluation metrics, with the benefit of the structure guidance from the intermediate frames. Furthermore, Fig. 6 shows qualitative comparison between video frame interpolation methods FILM [36] and MaskINT. Even when confronted with significant motion between two frames, our approach successfully reconstructs the original video, maintaining consistent motion through the aid of structural guidance. In contrast, FILM introduces undesirable artifacts, including disorted background, duplicated cat hands, and the absence of a camel’s head, etc. The major reason is that current VFI models mainly focus on generating slow-motion effects and enhancing frame rate, making them less effective in handling frames with large motions, which usually requires a better semantic understanding. Additionally, the absence of structural guidance poses a challenge for these VFI methods in accurately aligning generated videos with the original motion.
@@
Although our model is trained with frame interpolation by default, MaskINT can seamlessly generalize to an arbitrary number of keyframes without finetuning. We assess the impact of varying the quantity of keyframes on the generation performance. As shown in the left part of Table 3, with an increase in the number of keyframes, the model exhibits an improvement in performance. Generally, with more information, performing frame interpolation is easier. However, simultaneously editing more keyframes requires longer time due to the global attention among them.
@@
We also explore the number of decoding steps of the masked generative transformers in the second stage. The right part of Table 3 shows that more decoding steps can bring slight improvement on the temporal consistency, but requires more time. Considering the trade-off between performance and efficiency, we chose $K=32$ steps by default in all experiments.
# keyframes T.C. P.C. Time 1 0.9690 0.2984 19s 2 0.9714 0.3038 22s 3 0.9721 0.3051 26s 4 0.9728 0.3069 29s 6 0.9737 0.3035 35s
#decoding step $K$ T.C. P.C. Time 16 0.9691 0.3038 15s 32 0.9714 0.3038 22s 64 0.9719 0.3040 33s 128 0.9720 0.3041 62s
Table 3: Ablation study on ShutterStock dataset of the number of keyframes and the number of decoding steps $K$. “T.C.” stands for “temporal consistency” and “P.C.” stands for “prompt consistency”.
@@ 6 Limitation and Future Work
One potential limitation of our approach is that it still necessitates the use of diffusion-based methods for keyframe editing in the initial stage, which can still be somewhat time-consuming. As for the future work, given that our approach disentangles video editing into two distinct stages, we intend to explore the integration of token-based methods, like Muse [8], for image editing. This endeavor aims to further enhance the efficiency of the initial stage.

### v2

@@ 4.3 Structure-Aware Frame Interpolation
Non-autoregressive masked generative transformers [20, 60] can effectively perform frame prediction and interpolation tasks. However, these works lack explicit control of structure, making it difficult to follow the motion of input videos. Thus, we explicitly introduce structural condition of each frame into the generation process. In detail, we tokenize both RGB frames $\{\mathbf{I}^{i}\}_{i=1}^{N}$ and structure maps $\{\mathbf{S}^{i}\}_{i=1}^{N}$ with an off-the-shelf 2D VQ tokenizer [16]. We utilize 2D VQ rather than 3D VQ [60] to accommodate diverse numbers of frames and frame rate. We denote tokens from RGB frames as $\mathbf{{}^{c}Z}=\{^{c}z_{i}\}_{i=1}^{N\times h\times w}$ (color token) and tokens from edge maps as $\mathbf{{}^{s}Z}=\{^{s}z_{i}\}_{i=1}^{N\times h\times w}$ (structure token), where ${}^{s}z_{i}$, ${}^{c}z_{i}\in\{1,2,...,M\}$ and $M$ is the codebook size. Subsequently, two distinct embedding layers $e^{C}(\cdot)$ and $e^{S}(\cdot)$ are learned to map tokens $\mathbf{{}^{c}Z}$ and $\mathbf{{}^{s}Z}$ into their respective embedding spaces. Learnable 2D spatial positional encoding $\mathbf{P^{S}}$ and temporal positional encoding $\mathbf{P^{T}}$ are also added [4]. Thus, the input to the following transformer layers can be formulated by $\mathbf{X}=e^{c}(\mathbf{{}^{c}Z})+e^{s}(\mathbf{{}^{s}Z})+\mathbf{P^{S}}+\mathbf{P^{T}}\in\mathbb{R}^{N\times h\times w\times c}$.
@@
Previous masked generative transformers [10, 60] employ the standard transformer with global attention [53]. Given that there is no substantial motion between consecutive frames, we adopt self-attention within a restricted window [20] to further mitigate computational overhead In detail, our approach involves two distinct stages of attention. First, we conduct spatial window attention, where attention is calculated within each frame of dimensions $1\times h\times w$. We then conduct spatial-temporal window attention, where attention is calculated within a tube of dimensions $N\times h_{w}\times w_{w}$, where $h_{w}$ and $w_{w}$ is the window size. Besides, to further reduce computation of transformer, we also add a shallow convolution layers to downsample $\mathbf{X}$ before the transformer encoder layers and an upsample layer at the end.
@@
We train MaskINT using video-only datasets. Denote the color token of $i^{th}$ RGB frame as $\mathbf{{}^{c}Z}^{i}=\{^{c}z^{i}\}_{i=1}^{h\times w}$. During the training time, we keep color tokens of the initial frame $\mathbf{{}^{c}Z}^{1}$ and last frame $\mathbf{{}^{c}Z}^{N}$, and randomly replace $[\gamma(r)\cdot(N-2)\cdot N]$ color tokens of intermediate frames with the [MASK] tokens. We denote this corrupted video color tokens as $\mathbf{\bar{{}^{c}Z}}=\{\mathbf{{}^{c}Z}^{1},\mathbf{\bar{{}^{c}Z}}^{2},...,\mathbf{\bar{{}^{c}Z}}^{N-1},\mathbf{{}^{c}Z}^{N}\}$. The structure-aware window-restricted transformer with parameters $\Theta$ is trained by
$\mathcal{L}_{MTM}=\mathop{\mathbb{E}}_{\{\mathbf{{}^{c}Z},\mathbf{{}^{s}Z}\}\in\mathbb{D}}[\sum_{\bar{{}^{c}z_{i}}=\texttt{[MASK]}}-\log p_{\Theta}(^{c}z_{i}|\mathbf{\bar{{}^{c}Z}},\mathbf{{}^{s}Z})].$ (4)
@@
During inference, MaskINT can seamlessly generalize to perform frame interpolation between the jointly edited frames, although it is only trained with regular videos. Specifically, we tokenize the the initial and last edited frames $\mathbf{{}^{*}I}^{1}$ and $\mathbf{{}^{*}I}^{N}$ from Stage 1 into color tokens $\mathbf{{}^{c}_{*}Z}^{1}$ and $\mathbf{{}^{c}_{*}Z}^{N}$, and initialize color tokens of all intermediate frames $\{\mathbf{\bar{{}^{c}Z}}^{2},...,\mathbf{\bar{{}^{c}Z}}^{N-1}\}$ with [MASK] tokens. We follow the iterative decoding in MaskGiT [10] with a total number of $K$ steps. At step $k$, we predict all color tokens in parallel and keep tokens with the highest confidence score.
@@ 5.1 Settings
We train our model with $100k$ videos from ShutterStock [1]. During training, we random select a video clip with $T=16$ frames and frame interval of $1,2,4$ from each video. All frames are resized to $384\times 672$. We utilize Retina-VQ [15] with $8$ downsample ratio, i.e., each frame has $48\times 84$ tokens. We employ Transformer-Base as our MaskINT and optimized it from scratch with the AdamW optimizer [33] for a duration of $100$ epochs. The initial learning rate is set to $41e-$ 4 and decayed with cosine schedule. We set the number of decoding step $K$ to $32$ and the temperature $t$ to $4.5$ in inference.
@@
We evaluate our method on $40$ selected object-centric videos [56] from the DAVIS dataset [40], as well as $30$ unseen videos from the ShutterStock [1], covering animals, vehicles, etc. For each video, we manually design $5$ edited prompts, including object editing, background changes and style transfers. Following previous works [62, 41, 19, 8], we evaluate the quality of the generated videos with the following metrics: 1) Prompt consistency, which calculates the CLIP [42] embedding similarity between text prompt and all video frames. 2) Temporal consistency, which calculates the CLIP embedding similarity of all pairs of consecutive frames. 3) Warpping-error [30], which computes the optical flow between consecutive frames using RAFT [50], and warps the edited frames to the next according it. 4) Long-term temporal consistency. Inspired by the global coherence metric in [52], we also calculate the warp error between all possible pairs of frames as the long-term consistency. As for efficiency, we report the running time for generating a 16-frame video clip on a single NVIDIA A6000 GPU.
@@ 5.2 Results
We apply ControlNet to each frame individually with the same noise as the baseline. We select methods that built upon T2I diffusion models for comparison, including Tune-A-Video [56], TokenFlow [19], Text-to-video zero [29], and ControlVideo [62]. Besides, we compare with VFI methods such as FILM [43] using the same edited keyframes.
@@
[…]
/×$\times 10^{-3}$ /×$\times 10^{-3}$ /×$\times 10^{-3}$ /×$\times 10^{-3}$
ControlNet per frame [61] 0.314 0.914 36.6 58.6 0.304 0.942 25.9 45.1 50s
Tune-a-Video [56] 0.299 0.966 20.4 48.2 0.292 0.979 13.4 34.9 20min
Text2Video-zero [29] 0.312 0.964 20.7 42.5 0.304 0.981 16.0 33.0 60s
ControlVideo-edge [62] 0.314 0.975 6.9 23.9 0.303 0.986 7.4 20.5 120s
TokenFlow [19] 0.317 0.977 7.0 19.6 0.313 0.987 5.4 15.8 150s
MaskINT (ours) 0.311 0.952 9.5 27.7 0.304 0.971 8.6 22.3 22s
Table 1: Quantitative comparisons. “T.C." stands for “temporal consistency", “P.C." stands for “prompt consistency", “W.E" stands for “warpping-error", and “L.C." stands for “long-term temporal consistency".
@@
Tab. 1 summarize the performance of these methods on both DAVIS and ShutterStock datasets. Notably, our method achieves comparable performance with diffusion methods, in terms of all four evaluation metrics, while brings a significant acceleration in processing speed. In detail, the test-time finetuning of TAV [56] is extremely time-consuming, making it a less practical solution. MaskINT is almost 5.5 times faster than ControlVideo [62], whose fully cross-frame attention is computationally extensive. Moreover, MaskINT is nearly 7 times faster than TokenFlow [19], whose DDIM inversion is time-consuming. The efficiency of our method is derived from a combination of a lightweight network and a reduced number of decoding steps with masked generative transformers.
@@
Fig. 1 and Fig. 3 show edited video samples using MaskINT. Our method is able to generate temporally consistent videos in accordance with text prompts, enabling diverse applications including stylization, background and foreground editing, and beyond. It is also effective in videos with complex motion, such as jumping and running, as well as scenarios with multiple subjects. Besides, although the structure-aware interpolation module is trained with realistic videos only, it still generalizes to diverse synthesized images. For example, we can successfully generate a video in the style of Van Gogh even in the absence of videos for this particular style in the training set.
Fig. 4 provides a comparison of MaskINT to other methods. Remarkably, diffusion-based methods with extended attention [56, 19, 29, 62] can maintain the overall appearance consistency, but sometimes result in inconsistent details. For example, TokenFlow [19] and Text2Video-Zero[29] exhibit noticeable artifacts in the leg region of the human subjects and ControlVideo [62] produces inconsistent hats. The potential explanation lies in the fact that these methods offer control over temporal consistency implicitly. Besides, FILM [43] generates videos that cannot follow the original motions. Our MaskINT consistently interpolates the intermediate frames based on the structure condition and even maintain better consistency in local regions.
@@
Given that the non-autoregressive pipeline generates all frames simultaneously, it’s challenging to edit an entire long video at once due to GPU memory limitation. Nevertheless, our framework can be extended to handle long videos by dividing them into short clips and progressively performing frame interpolation within each segment. For instance, given a video with $60$ frames, we select the $1^{\text{st}}$, $16^{\text{th}}$, $31^{\text{st}}$, $46^{\text{th}}$, and $60^{\text{th}}$ frames as keyframes. We jointly edit these selected $5$ frames together and then perform structure-aware frame interpolation within each pair of consecutive keyframes. As shown in Fig. 5, with this design, our method can still generate consistent long videos. Moreover, in this proposed extension, the generation of later frames is dissociated from that of early frames, which differs from the autoregressive generation pipeline in Phenaki [54]. Consequently, even if some early frames encounter difficulties, the generation of later frames can still proceed successfully.
@@ 5.3 Ablation Studies
Video Frame Interpolations.
In this ablation study, we evaluate the performance of the structure-aware interpolation module. In detail, we perform frame interpolation using original keyframes, and compare the interpolated frames with the original intermediate video frames. We use same testing samples from DAVIS and Shutterstock datasets, employing peak signal-to-noise ratio (PSNR), learned perceptual image patch similarity (LPIPS), and structured similarity (SSIM) as the evaluation metrics. We benchmark our method against two state-of-the-art VFI methods, namely FILM [43] and RIFE [26]. We also showcase the performance of applying VQ-GAN [16] to all video frames, serving as an upper bound for our method.
As in Tab. 2, our method significantly outperforms VFI methods on all evaluation metrics, with the benefit of the structure guidance from the intermediate frames. Furthermore, Fig. 6 shows qualitative comparison between FILM [43] and MaskINT. Even when confronted with significant motion between two frames, our method successfully reconstructs the original video, maintaining consistent motion through the aid of structural guidance. In contrast, FILM introduces undesirable artifacts, including disorted background, duplicated cat hands, and the absence of a camel’s head, etc. The major reason is that current VFI models mainly focus on generating slow-motion effects and enhancing frame rate, making them less effective in handling frames with complex motions, which usually requires a better semantic understanding. Besides, the absence of structural guidance poses a challenge for them in accurately aligning generated videos with the original motion.
@@
Although our model is trained with frame interpolation by default, MaskINT can seamlessly generalize to an arbitrary number of keyframes without finetuning. We assess the impact of varying the quantity of keyframes on the generation performance. As shown in the left part of Table 3, with an increase in the number of keyframes, the model exhibits an improvement in performance. Generally, with more information, performing frame interpolation is easier. However, simultaneously editing more frames requires longer time due to the global attention among them.
@@
We also explore the number of decoding steps $K$ for the masked generative transformers in the second stage. The right part of Tab. 3 shows that more decoding steps can bring slight improvement on the temporal consistency, but requires more time. Considering the trade-off between performance and efficiency, we chose $K=32$ steps by default in all experiments.
# keyframes T.C. P.C. Time
1 0.9690 0.2984 19s
2 0.9714 0.3038 22s
3 0.9721 0.3051 26s
4 0.9728 0.3069 29s
6 0.9737 0.3035 35s
#decoding step $K$ T.C. P.C. Time
16 0.9691 0.3038 15s
32 0.9714 0.3038 22s
64 0.9719 0.3040 33s
128 0.9720 0.3041 62s
Table 3: Ablation study on the number of keyframes and the number of decoding steps $K$. “T.C." stands for “temporal consistency" and “P.C." stands for “prompt consistency".
@@ 6 Limitation and Future Work
One limitation of MaskINT is that it can only perform structure-preserving video editing, such as altering style or appearance. Thus, it cannot handle edits that require structural changes, for example, change a dog to a horse, a limitation shared with TokenFlow [19]. It also requires that no new objects should appear in the intermediate frames. Besides, the performance of MaskINT highly relies on the tedious image-editing model and structure detector. When these models fail, the structure-aware interpolation becomes meaningless, resulting in artifacts. In the future, given that our approach disentangles video editing into two distinct stages, we intend to explore the integration of token-based methods, like Muse [11], for image editing. This endeavor aims to further enhance the efficiency of the initial stage.